A man borrows 10,000 from a loan firm. the rate of simple interest is 15% but the interest is to be deducted from the loan at the time the money is borrowed. at the end of one year he has to pay back 10000. what is the actual rate of interest ?

1
Expert's answer
2022-02-18T08:43:17-0500

Actual Rate of Interest

Borrowed amount "=10,000"

"r=15\%"

"Interest\ at\ borrowing=(0.15\times10,000)=1500"

Received amount "=10,000-1500=8500"

Pay amount "=10,000"

Difference between paid and received amount "=(10,000-8500)=1500"

"\frac{(1500-10,000)}{10,000}\times100=15\%"= Interest at Paying

"Actual\ Interest =" "Interest\ at\ borrowing + Interest\ at\ Paying"

"Actual\ interest=15\%+15\%=\bold{30\%}"